In a sample of 300 undergraduates, 90 males and 65 females were in the College of Arts and Sciences, 45 males and 36 females were in the College of Business, and 30 males and 34 females were in the College of Education. If one student is randomly selected, find the probability that the student is

a) A female in the College of Business

b) A male in the College of Education

c) A female

d) A student in the College of A&S

e) A male in the College of Business

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

36/300=12% female in business college
30/300=10% male in ed. college
135/300=45% a female
155/300=51 2/3% in a&s
45/300=15% a male in bus. college

Answer by stanbon(48568)

Form a contingency table with
rows for "Arts & Science", Business, Education.
columns for male and female.
----------
calculate the row and the column totals
-----------
Comment: Your questions are somewhat ambiguous.
For example Probability of"a female in the
College Of Arts and Science" could be 65/165
If you limit the choices to the number in the
College of Arts and Science: or the answer
could be 65/300 if you are only limited by
the number of students. I'm going to take
the latter approach and say the answer is 65/300
for question "a".
------------
b) 30/300
c) 135/300
d) 155/300
e) 45/300
-------------
If you construct the contingency table you will
see the ambiguity I see in your questions.
Cheers,
Stan H.